A 0.15kg tumor is treated with α

α particles. There are 2×10

12

2×1012 particles reaching the tumor every second, each having energy 3×10

−15

3×10−15J. What is the absorbed dose given to the tumor



1
Expert's answer
2021-11-25T16:24:09-0500

"m=0.15 \; kg \\

n=2 \times 10^{12} \; particles \\

E = 3 \times 10^{-15} \;J"

The absorbed dose "= \frac{nE}{m}"

"= \frac{2 \times 10^{12} \times 3 \times 10^{-15}}{0.15} \\

= 40 \times 10^{-3} \\

= 0.04 \;J/kg \cdot s"